A body is projected upward with a velocity of 30m/s. Find the time when the velocity will be 5m/s


1
Expert's answer
2021-06-29T10:00:38-0400

Under the influence of gravity, the object decelerates at the rate of 9.8 m/s2. So, the time is


"t=\frac{v_2-v_1}{g}=\frac{5-30}{-9.8}=2.55\text{ s}."